# Bounds reach of a future circular collider on the parameters of a minimal Baryon-Lepton symmetric model

### v2

@@ Abstract
This paper investigates the physics potential of the proton–proton Future Circular Collider (FCC-$h$) in the search for a new heavy gauge boson, $Z^{\prime}$, within the framework of the minimal Baryon-Lepton (BLB-L) symmetric model. We examine the exclusion limits on the mass of the $Z^{\prime}$ boson for various sets of coupling constants, utilizing leptonic decays, specifically $\ell^{+}\ell^{-}$. Using the DARKCAST framework, we compare the parameter constraints derived from previous experiments with the expected reach of the FCC-$h$, providing insights into the collider’s sensitivity and potential to probe new physics beyond the Standard Model.
@@ 1 Introduction
Despite decades of experimental effort, the direct discovery of dark matter, new gauge bosons such as the $Z^{\prime}$, or any particle of BSM has remained elusive. The Large Hadron Collider (LHC) has played a central role in testing the Standard Model and searching for new physics at the TeV scale. While it has not provided evidence for new particles beyond the Higgs boson, the LHC has established stringent limits on many extensions of the SM, including searches for additional neutral gauge bosons. To further extend the energy frontier and to explore parameter regions that remain inaccessible, future collider projects have been proposed. Among these, the hadronic Future Circular Collider (FCC-$h$) is designed to reach a center-of-mass energy of about 100 TeV with an integrated luminosity of 30 ab-1. Such capabilities would allow precision measurements of SM processes at unprecedented scales and dramatically improve the discovery potential for heavy new particles such as the $Z^{\prime}$. In parallel, the proposed International Linear Collider (ILC) is designed to operate as a high-precision $e^{+}e^{-}$ facility with center-of-mass energies up to $1~\text{TeV}$. In the context of the BLB\!-\!L model, these precision measurements are particularly relevant, as the anomaly-free realization of the gauge symmetry implies an upper theoretical limit on the parameter ratio $M_{Z^{\prime}}/g^{\prime}\gtrsim 137.2~\text{TeV}$ [Das:2021esm, Babich:2010zz, Pankov:2017dkv]. The ILC therefore provides a complementary probe of the model by constraining the coupling structure even when the $Z^{\prime}$ mass lies well above its direct kinematic reach. Other proposed facilities, including the High-Luminosity LHC (HL-LHC) [Apollinari:2015wtw, LHeC:2020van], the High-Energy LHC (HE-LHC) [FCC:2018bvk], and the Circular Electron–Positron Collider (CEPC) [CEPC-SPPCStudyGroup:2015csa, CEPCStudyGroup:2023quu], provide important complementary opportunities. Nevertheless, among these, the FCC-$h$ offers the broadest reach for probing uncharted regions of parameter space in both the SM and its extensions [FCC:2018evy, FCC:2018vvp, FCC:2018byv, Armesto:2014iaa].
The following sections are organized as follows. In the next chapter, we provide a detailed discussion of the BLB\!-\!L model, outlining its theoretical structure and phenomenological features relevant for collider studies. The subsequent chapter then presents the simulation setup adopted in this work, including the choice of parameters, event generation, and analysis strategy used to evaluate the sensitivity of the FCC-$h$ to the $Z^{\prime}$.
@@ 3.2 Detector Response
The detector response has been simulated via the DELPHES[deFavereau:2013fsa] software package where default FCC-$h$ with pile-up card of DELPHES version 3.5.0 was used. Following a collision event, which involves parton showering, hadronization, and particle decays, DELPHES initiates its process by ensuring the propagation of long-lived particles through the tracking volume. This volume is situated within a uniform axial magnetic field, aligned parallel to the beam direction. The resolution $\sigma$ on the track transverse momentum $p_{T}$, is described as,
@@ 4 Results and discussion
Figure 3 presents the limits on the $Z^{\prime}$ coupling and mass, incorporating our new simulation results alongside the existing experimental constraints. The exclusion regions shown in the plot are obtained using DARKCAST [Ilten:2018crw], which we employed to reinterpret data from previous dark-photon experiments in the context of the BLB\!-\!L model. In particular, the software allows for a model-independent translation of published limits on the kinetic-mixing parameter $\epsilon$ of a dark photon into corresponding limits on the gauge coupling $g^{\prime}$ of a BLB\!-\!L boson by rescaling the production and decay rates. This procedure provides a unified framework for comparing dark-photon searches with other $U(1)_{X}$ extensions of the Standard Model. Each shaded region represents a parameter space excluded by previous experiments under different physical assumptions. The low-mass domain ($m_{Z^{\prime}}\lesssim 10~\text{GeV}$) is primarily constrained by fixed-target and beam-dump experiments, whereas the intermediate and high-mass ranges are probed by collider and precision measurements. The dark green, red, and brown shaded regions at the lower left correspond to classic beam-dump and neutrino experiments such as E137, E141, PS191, CHARM, and NOMAD, which exclude extremely small couplings ($g^{\prime}\lesssim 10^{-7}$) for sub-GeV masses. Experiments such as APEX, A1, KLOE, NA48, and BaBar probe visible decays at low-mass region, excluding couplings in the range $10^{-4}\lesssim g^{\prime}\lesssim 10^{-2}$. At higher masses ($m_{Z^{\prime}}\gtrsim 10~\text{GeV}$), collider experiments dominate the sensitivity. The light-blue and orange shaded areas represent exclusion regions obtained by LHCb, CMS, and BaBar, which probe the dilepton final states through resonance searches. The CMS limits, in particular, extend up to a few TeV, excluding couplings above $g^{\prime}\sim 10^{-2}$. The magenta region labeled FCC shows the new sensitivity derived in this work using $30~\text{ab}^{-1}$ of simulated FCC-$h$ data. The projected exclusion demonstrates a substantial improvement over existing collider bounds, extending the reachable mass scale up to $\mathcal{O}(10^{4})~\text{GeV}$ while probing couplings down to $g^{\prime}\sim 10^{-2}$. This level of reach reflects the combined effect of the extremely high collision energy and integrated luminosity envisioned for the FCC.
@@ 5 Summary and Conclusions
In this work, we investigated the sensitivity of the Future Circular Collider (FCC-$h$) to a heavy $Z^{\prime}$ boson in the leptonic ($e^{+}e^{-},\,\mu^{+}\mu^{-}$) final states. The analysis focused exclusively on the dominant Standard Model Drell–Yan background, which rapidly decreases with increasing invariant mass, providing an almost background-free region in the multi-TeV regime. Using simulated FCC data corresponding to an integrated luminosity of $30~\text{ab}^{-1}$, we derived 95% CL upper limits on the production cross section $\sigma(pp\to Z^{\prime})\times\mathrm{BR}(Z^{\prime}\to\ell^{+}\ell^{-})$ for various coupling constants. For the largest coupling considered ($g^{\prime}=0.3$), the analysis excludes $Z^{\prime}$ masses up to approximately $45~\text{TeV}$, demonstrating the exceptional reach of a 100 TeV collider in probing physics beyond the Standard Model.
@@
In summary, this study demonstrates that the FCC-$h$, with its unprecedented center-of-mass energy and luminosity, provides a powerful environment to explore heavy neutral gauge bosons predicted by BLB\!-\!L and related $U(1)_{X}$ extensions. The derived limits represent a substantial improvement over existing bounds and establish a strong case for future experimental efforts to test the presence of new gauge interactions at the multi-TeV scale.

### v4

@@ Abstract
This paper investigates the physics potential of the proton–proton Future Circular Collider (FCC-hh) in the search for a new heavy gauge boson, $Z^{\prime}$, within the framework of the minimal Baryon-Lepton (BLB-L) symmetric model. We examine the exclusion limits on the mass of the $Z^{\prime}$ boson for various sets of coupling constants, utilizing leptonic decays, specifically $\ell^{+}\ell^{-}$. Using the DARKCAST framework, we compare the parameter constraints derived from previous experiments with the expected reach of the FCC-hh, providing insights into the collider’s sensitivity and potential to probe new physics beyond the Standard Model.
@@ 1 Introduction
Despite decades of experimental effort, the direct discovery of dark matter, new gauge bosons such as the $Z^{\prime}$, or any particle of BSM has remained elusive. The Large Hadron Collider (LHC) has played a central role in testing the Standard Model and searching for new physics at the TeV scale. While it has not provided evidence for new particles beyond the Higgs boson, the LHC has established stringent limits on many extensions of the SM, including searches for additional neutral gauge bosons. To further extend the energy frontier and to explore parameter regions that remain inaccessible, future collider projects have been proposed. Among these, the hadronic Future Circular Collider (FCC-hh) is designed to reach a center-of-mass energy of about 100 TeV with an integrated luminosity of 30 ab-1. Such capabilities would allow precision measurements of SM processes at unprecedented scales and dramatically improve the discovery potential for heavy new particles such as the $Z^{\prime}$. In parallel, the proposed International Linear Collider (ILC) is designed to operate as a high-precision $e^{+}e^{-}$ facility with center-of-mass energies up to $1~\text{TeV}$. In the context of the BLB\!-\!L model, these precision measurements are particularly relevant, as the anomaly-free realization of the gauge symmetry implies an upper theoretical limit on the parameter ratio $M_{Z^{\prime}}/g^{\prime}\gtrsim 137.2~\text{TeV}$ [Das:2021esm, Babich:2010zz, Pankov:2017dkv]. The ILC therefore provides a complementary probe of the model by constraining the coupling structure even when the $Z^{\prime}$ mass lies well above its direct kinematic reach. Other proposed facilities, including the High-Luminosity LHC (HL-LHC) [Apollinari:2015wtw, LHeC:2020van], the High-Energy LHC (HE-LHC) [FCC:2018bvk], and the Circular Electron–Positron Collider (CEPC) [CEPC-SPPCStudyGroup:2015csa, CEPCStudyGroup:2023quu], provide important complementary opportunities. Nevertheless, among these, the FCC-hh offers the broadest reach for probing uncharted regions of parameter space in both the SM and its extensions [FCC:2018evy, FCC:2018vvp, FCC:2018byv, Armesto:2014iaa].
The following sections are organized as follows. In the next chapter, we provide a detailed discussion of the BLB\!-\!L model, outlining its theoretical structure and phenomenological features relevant for collider studies. The subsequent chapter then presents the simulation setup adopted in this work, including the choice of parameters, event generation, and analysis strategy used to evaluate the sensitivity of the FCC-hh to the $Z^{\prime}$.
@@ 3.2 Detector Response
The detector response has been simulated via the DELPHES[deFavereau:2013fsa] software package where default FCC-hh with pile-up card of DELPHES version 3.5.0 was used. Following a collision event, which involves parton showering, hadronization, and particle decays, DELPHES initiates its process by ensuring the propagation of long-lived particles through the tracking volume. This volume is situated within a uniform axial magnetic field, aligned parallel to the beam direction. The resolution $\sigma$ on the track transverse momentum $p_{T}$, is described as,
@@ 4 Results and discussion
Figure 3 presents the limits on the $Z^{\prime}$ coupling and mass, incorporating our new simulation results alongside the existing experimental constraints. The exclusion regions shown in the plot are obtained using DARKCAST [Ilten:2018crw], which we employed to reinterpret data from previous dark-photon experiments in the context of the BLB\!-\!L model. In particular, the software allows for a model-independent translation of published limits on the kinetic-mixing parameter $\epsilon$ of a dark photon into corresponding limits on the gauge coupling $g^{\prime}$ of a BLB\!-\!L boson by rescaling the production and decay rates. This procedure provides a unified framework for comparing dark-photon searches with other $U(1)_{X}$ extensions of the Standard Model. Each shaded region represents a parameter space excluded by previous experiments under different physical assumptions. The low-mass domain ($m_{Z^{\prime}}\lesssim 10~\text{GeV}$) is primarily constrained by fixed-target and beam-dump experiments, whereas the intermediate and high-mass ranges are probed by collider and precision measurements. The dark green, red, and brown shaded regions at the lower left correspond to classic beam-dump and neutrino experiments such as E137, E141, PS191, CHARM, and NOMAD, which exclude extremely small couplings ($g^{\prime}\lesssim 10^{-7}$) for sub-GeV masses. Experiments such as APEX, A1, KLOE, NA48, and BaBar probe visible decays at low-mass region, excluding couplings in the range $10^{-4}\lesssim g^{\prime}\lesssim 10^{-2}$. At higher masses ($m_{Z^{\prime}}\gtrsim 10~\text{GeV}$), collider experiments dominate the sensitivity. The light-blue and orange shaded areas represent exclusion regions obtained by LHCb, CMS, and BaBar, which probe the dilepton final states through resonance searches. The CMS limits, in particular, extend up to a few TeV, excluding couplings above $g^{\prime}\sim 10^{-2}$. The magenta region labeled FCC shows the new sensitivity derived in this work using $30~\text{ab}^{-1}$ of simulated FCC-hh data. The projected exclusion demonstrates a substantial improvement over existing collider bounds, extending the reachable mass scale up to $\mathcal{O}(10^{4})~\text{GeV}$ while probing couplings down to $g^{\prime}\sim 10^{-2}$. This level of reach reflects the combined effect of the extremely high collision energy and integrated luminosity envisioned for the FCC.
@@ 5 Summary and Conclusions
Figure 3: Limits of the coupling and mass of $Z^{\prime}$ with respect to previous experiments and the extended limits of this bounds determined from the simulation.
In this work, we investigated the sensitivity of the Future Circular Collider (FCC-hh) to a heavy $Z^{\prime}$ boson in the leptonic ($e^{+}e^{-},\,\mu^{+}\mu^{-}$) final states. The analysis focused exclusively on the dominant Standard Model Drell–Yan background, which rapidly decreases with increasing invariant mass, providing an almost background-free region in the multi-TeV regime. Using simulated FCC data corresponding to an integrated luminosity of $30~\text{ab}^{-1}$, we derived 95% CL upper limits on the production cross section $\sigma(pp\to Z^{\prime})\times\mathrm{BR}(Z^{\prime}\to\ell^{+}\ell^{-})$ for various coupling constants. For the largest coupling considered ($g^{\prime}=0.3$), the analysis excludes $Z^{\prime}$ masses up to approximately $45~\text{TeV}$, demonstrating the exceptional reach of a 100 TeV collider in probing physics beyond the Standard Model.
@@
In summary, this study demonstrates that the FCC-hh, with its unprecedented center-of-mass energy and luminosity, provides a powerful environment to explore heavy neutral gauge bosons predicted by BLB\!-\!L and related $U(1)_{X}$ extensions. The derived limits represent a substantial improvement over existing bounds and establish a strong case for future experimental efforts to test the presence of new gauge interactions at the multi-TeV scale.